If a student owns a computer system worth $4,000, how much would insurance cost to cover the system for 146 days if the insurance rate it 1.5% per year?
Answer by stanbon(75887)

If a student owns a computer system worth $4,000, how much would insurance cost to cover the system for 146 days if the insurance rate it 1.5% per year?
------------------------------
Ans: (146/365)(0.015)4000 = 0.4*60 = $24
====================================================
Cheers,
Stan H.
==============